### v1

# Semi-supervised and unsupervised learning for health indicator extraction from guided waves in aerospace composite structures

Abstract
Health indicators (HIs) are central to diagnosing and prognosing the condition of aerospace composite structures, enabling efficient maintenance and operational safety. However, extracting reliable HIs remains challenging due to variability in material properties, stochastic damage evolution, and diverse damage modes. Manufacturing defects (e.g., disbonds) and in‑service incidents (e.g., bird strikes) further complicate this process. This study presents a comprehensive data‑driven framework that learns HIs via two learning approaches integrated with multi‑domain signal processing. Because ground‑truth HIs are unavailable, a semi-supervised and an unsupervised approach are proposed: (i) a diversity deep semi-supervised anomaly detection (Diversity-DeepSAD) approach augmented with continuous auxiliary labels used as hypothetical damage proxies, which overcomes the limitation of prior binary labels that only distinguish healthy and failed states while neglecting intermediate degradation, and (ii) a degradation-trend-constrained variational autoencoder (DTC-VAE), in which the monotonicity criterion is embedded via an explicit trend constraint. Guided waves with multiple excitation frequencies are used to monitor single‑stiffener composite structures under fatigue loading. Time, frequency, and time–frequency representations are explored, and per‑frequency HIs are fused via unsupervised ensemble learning to mitigate frequency dependence and reduce variance. Using fast Fourier transform features, the augmented Diversity-DeepSAD model achieved 81.6% performance, while DTC-VAE delivered the most consistent HIs with 92.3% performance, outperforming existing baselines. This work contributes: (i) a Diversity-DeepSAD variant for intermediate degradation modeling via continuous labels, (ii) the first application of Diversity-DeepSAD and DTC-VAE to guided wave monitoring of aerospace composites, and (iii) a framework for history-free HI estimation. These advancements provide reliable HIs, reducing maintenance costs and enhancing safety in aerospace structures.
keywords:
Prognostics and health management , Health indicator , Constrained variational autoencoder , Semi-supervised learning , Aeronautical composite structures
[…]
[delft]organization=Center of Excellence in Artificial Intelligence for Structures, Prognostics & Health Management, Aerospace Engineering Faculty, Delft University of Technology, addressline=Kluyverweg 1, city=Delft, postcode=2629 HS, country=the Netherlands
List of Symbols $c$ Hypersphere centre $D$ Depth of network $F_{all}$ Fitness function $F_{test}$ Test fitness function $f$ Frequency $H$ Number of statistical features $h$ Statistical feature $i,j$ Timestep $k$ Prognostic criteria weighting $\mathcal{L}$ Loss function $M$ Total number of samples MoMo Monotonicity $m$ Sample number NN Number of measurements (timesteps) $N_{f}$ Number of frequencies $n$ Number of labeled samples $P_{\theta}$ Decoder probability distribution PrPr Prognosability $Q_{\phi}$ Encoder probability distribution $r$ Rate of degradation $S$ Frequency domain statistical feature $s(f)$ Frequency domain signal $t_{i}$ Time at timestep $i$ TrTr Trendability $u$ Number of unlabeled samples $W_{l}$ Network weights of layer $l$ $X$ Time domain statistical feature $x(i)$ Time domain signal $x_{i}^{m,f}$ Features at timestep $i$ of specimen $m$ with measurement frequency $f$ $\hat{x}$ Reconstructed features $y_{i}^{m,f}$ Health indicator at timestep $i$ of specimen $m$ with measurement frequency $f$ $\tilde{y}$ Training label $\bar{y}$ Mean health indicator across folds $z$ Variational autoencoder latent variables $\alpha$ Kullback-Leibler divergence weighting $\beta$ Reconstruction loss weighting $\gamma$ Monotonicity constraint loss weighting $\epsilon$ Gaussian noise $\zeta_{k}$ $k$th eigenvalue of Gram matrix $\eta$ Labelled parameter loss weighting $\theta$ Neural network function $\lambda$ Diversity loss weighting $\mu$ Mean $\nu$ L2 regularisation weighting $\sigma$ Standard deviation $T$ Test specimen $\tau$ Training specimen $\phi$ Encoder parameters $\psi$ Decoder parameters $\omega_{f}$ Weight for frequency $f$
1 Introduction
The health of aerospace structures must be continuously monitored to ensure their safety and reliability. An effective monitoring and diagnostic system should detect damage early enough and issue warnings before it escalates into a dangerous situation [1]. Even earlier diagnostics using damage trend prediction, known as prognostics, can save significant time and cost by enabling timely repairs before a structure reaches an irreparable state [2, 3], in addition to improving reliability [4, 5, 6]. In this context, developing a health indicator (HI) is essential for both diagnostics and prognostics [7, 8, 9, 10]. A HI, being a value which quantifies the health of a structure, serves as a critical bridge between these two aspects towards condition-based maintenance (CBM) [11, 12]. However, deriving a truly comprehensive and accurate HI is exceptionally challenging for most objects, especially engineering systems, due to their inherent complexity [13]. The comprehensiveness of a HI refers to its ability to account for all potential types of damage that could affect a system’s health [14, 15, 13]. Tracking and quantifying all forms of degradation, even in simple objects made of isotropic materials, is often infeasible. Many degradation processes occur internally, beyond the reach of direct measurement. Despite the fact that sensors may monitor the effects of different types of damage on the sensed signals, meticulously identifying the damage remains challenging due to numerous known and unknown factors, which depend on the object under monitoring, type of sensors, and structural health monitoring (SHM) techniques employed. This complexity is further exacerbated in the case of advanced systems and materials [14, 16]. In summary, claiming to have completely true HI values for any object, especially a complex one like an engineering system, is an overstatement [13, 17].
While a perfectly accurate HI is impractical, certain principles about HIs hold universally true. For example, the end of life (EoL) represents the complete loss of health in any system. Similarly, a system’s health inevitably decreases over time unless maintenance or self-healing mechanisms are applied. These principles form the foundation for evaluating a HI using key metrics. Prognostability (Pr), for instance, ensures that the failure state at EoL is represented by the same HI value across different systems [18, 19]. Monotonicity (Mo) emphasises the consistency of health degradation over time [20, 18, 21, 22]. Another desirable metric is trendability (Tr), which assesses whether the degradation trends of similar systems are consistent. Although uncertainties in manufacturing, loading conditions, and environmental factors may introduce variability, the general trends for similar systems should not diverge significantly. Therefore, ideal HIs for a group of similar systems should satisfy Mo, Pr, and Tr [19, 13, 23, 24] (considering all in a fitness score), with the first two being fact-based and the third being aspirational but still valuable.
As already mentioned, developing or identifying HIs is a challenging task due to the inherent complexity and uncertainty of material degradation processes. This difficulty is amplified in anisotropic materials like composites [25] which are widely used in aerospace applications thanks to their high strength-to-weight ratio and ability to withstand directional stresses [26]. Unlike isotropic materials, composite materials exhibit more complex failure modes arising from their anisotropic properties and the potential interactions between different types of damage [27, 28, 15]. These unique characteristics make designing effective HIs for composite structures particularly intricate [14, 23]. Additionally, uncertainties arising during manufacturing (e.g., disbond defects) or operation (e.g., bird strikes) exacerbate this complexity [29, 30]. To enable in situ monitoring and account for these uncertainties, SHM techniques are essential. An effective SHM technique must be capable of covering most structural areas, particularly critical regions. Among these techniques, guided wave (GW)-based SHM stands out as a practical approach, where multiple sensors are strategically attached to critical regions to monitor the structure’s health. GWs are ultrasonic waves bounded by a structure’s material [30] that produce data from which the location and severity of damage to a structure can be determined [27, 31, 32, 33]. This enables SHM in otherwise inaccessible regions of complex composite structures [34]. As GW and most other SHM techniques rely on a network of sensors, model-based approaches face limitations in processing the large volume of complex, high-dimensional sensed signals. Furthermore, noise, multimodality, and the influence of environmental factors make it more challenging to generate HIs from raw GW data [15, 35, 36]. Consequently, data-driven and artificial intelligence (AI) approaches offer greater potential for handling such intricate datasets. Due to the unknown true health state of any engineering system, supervised models face limitations in constructing HIs. To address this challenge, semi-supervised or unsupervised AI models can be utilised for data fusion, enabling the generation of effective HIs [37, 38].
Semi-supervised models begin with concepts developed from supervised methods. For example, the support vector machine (SVM), introduced by Cortes and Vapnik [39], is a fundamental supervised classification algorithm. However, its binary classification ability falls short of generating meaningful HIs, as it ignores the complex gradient between healthy and failure states. Progress toward HI generation came with support vector data description (SVDD) [40] and its deep learning extension, Deep SVDD, proposed by Kim et al. in 2015 [41], which captures more intricate behaviors. For SHM, a semi-supervised model is preferred to effectively handle unlabeled data. Deep semi-supervised anomaly detection (DeepSAD), proposed by Ruff et al. in 2019 [42], extends Deep SVDD by incorporating unlabeled data points. While DeepSAD has proven effective in various non-SHM contexts, its application to HI generation poses challenges. For instance, Han et al. [43] used DeepSAD to detect anomalies in vehicle emissions, by setting a threshold anomaly score. They evaluated the performance of models by the area under the receiver operating characteristic curve, thus without considering the distribution of scores on either side of the threshold. This approach, although beneficial for anomaly detection, is inadequate for HI extraction due to the complete neglect of prognostic criteria. Similarly, DeepSAD has been applied with consideration only to the classification of anomalies in hydraulic systems [44]. Other approaches have sought to enhance semi-supervised anomaly detection. Gao et al. [45] developed ConNet, a robust deep anomaly detection model for sparsely labelled data. However, ConNet only considers positive, anomalous, labeled samples, which is unsuitable for HI generation as the healthy state is not considered.
Recent advancements have specifically targeted HI generation. Frusque et al. [46] applied DeepSAD to HI generation for rotating machinery, and it was observed that the objective function considers only the norm of the embedding, therefore producing low-rank results with some dimensions not fully utilised. To diversify the representation of degradation trends, they incorporated a diversity term in the loss function, creating Diversity-DeepSAD. For the PHME2010 milling dataset, Diversity-DeepSAD achieved promising results, with performances of 99% for Mo, 99% for Tr, and 94% for Pr, indicating room for improvement in Pr. However, when applied to thermal spray coating monitoring data, the model’s performance was lower, with 91% for Mo and 72% for Pr, while Tr was not reported. Based on the displayed HIs, the Tr score appears low. Notably, the coefficient of variation (CV), useful for comparing variability across datasets, was 24% for the thermal spray coating dataset. In contrast, aeronautical structures are significantly more challenging, with CV values of 51% for GW data from five single-stiffener composite panels [15] and 87% for acoustic emission data from 12 panels [23, 24]. Despite its potential, Diversity-DeepSAD has not yet been applied to aeronautical structures, which present unique challenges for HI construction and remaining useful life (RUL) prediction.
Moradi [13] developed various semi-supervised frameworks for single-stiffener composite structures. One approach utilised a multi-layer long short-term memory (LSTM) network, employing an intrinsically semi-supervised inductive learning technique, trained on acoustic emission-based time and frequency domain features [23]. This method achieved a fitness score of 93% under a rigorous leave-one-out cross-validation (LOOCV) process. However, the fine-tuning process to optimise the number of epochs should be done using only the validation specimen (unit), and another unit, different from the training/validation units, should be investigated to confirm the generalisability of this fine-tuned hyperparameter—something that was not considered in their study. To address this, Moradi et al. [24] introduced CEEMDAN-driven semi-supervised ensemble deep learning (CEEMDAN-SSEDL), which combined complete ensemble empirical mode decomposition with adaptive noise (CEEMDAN) for feature extraction and ensemble learning techniques to reduce randomness. By incorporating bidirectional LSTM (BiLSTM) layers during ensemble learning, this approach achieved a fitness score of 91.3%, while ensuring separate test units were considered. Moradi et al.’s [23] semi-supervised inductive learning approach was later applied by Frusque et al. [46] to the PHME2010 milling dataset, achieving a fitness score of 96%. However, the details of the architecture and the hyperparameter fine-tuning process were not reported. Recognising the limitations of historical-dependent models, Moradi et al. [37, 13] noted that such models often perform poorly in the absence of historical data. To overcome this, they introduced the Hilbert transform semi-supervised convolutional neural network (HT-SSCNN), designed to generate HIs from historical-independent GW data. Leveraging the GW monitoring technique, this approach achieved a 93% fitness score. Inspired by this, the methods in this paper are also designed to be historical-independent, relying solely on data from the current state of the structure. Given the demonstrated capabilities of anomaly detection algorithms, this study seeks to extend the application of Diversity-DeepSAD to the SHM of aerospace structures.
Previous applications of DeepSAD, however, have been limited to binary auxiliary labels distinguishing only healthy and failed states. This binary formulation neglects the intermediate degradation states that make up the majority of a structure’s lifetime, and in composites often leaves only a single label at failure, severely restricting the usefulness of labels. To address this gap, we propose the use of continuous auxiliary labels interpolated across the degradation process, providing richer supervision and smoother, more prognostic HIs.
Unsupervised models, by design, rely on feedback from the model output to learn without labeled data. A common technique is the autoencoder (AE), which uses a neural network to encode input data into a reduced-dimensional latent space and then reconstruct the input from this representation. Yang et al. [47] generated HIs for rotating machinery using sparse AEs, evaluating them based on the Mann-Kendall (MK) monotonicity metric. Lin and Tao [48] extended this approach by employing an ensemble of stacked AEs with linear targets, while Xu and Wang [49] enhanced stacked AEs with noise reduction using an exponential weight moving average model. In 2023, Xu et al. [50] implemented a deep convolutional AE and Mao et al. [51] applied tensor representation. These studies, however, focused on rotating machinery, where the vibrational behavior is more self-exposing and sensitive to damage. In contrast, aerospace structures do not exhibit such clear damage effects on the monitored signals, making the application of these methods more challenging. Variational autoencoder (VAE) architecture is similar to AE, but makes use of a probabilistic latent variable distribution in order to increase generability to data beyond the training set. In 2019, Ping et al. [52] utilised a VAE with a logarithmic normal distribution to generate HIs for turbofan engines (CMAPSS simulated dataset, which is simpler compared to experimental dataset [23, 53]). In 2020, Hemmer et al. [54] applied a VAE to rotating machinery at discrete damage levels. While generally increasing with damage level, the HIs do not conform to the prognostic criteria. Mitigating this, in 2022, Qin et al. [55] introduced a monotonic degradation constraint in the loss function, resulting in the degradation-trend-constrained VAE (DTC-VAE), which generates HIs that meet the Mo metric. Guo et al. [56] proposed a multiscale convolutional AE network for rotating machinery. Further studies in 2024 applied developments of VAE to the generation of HIs for bearings from vibrational data [57, 58, 59]. However, the application of VAEs for HI generation in aerospace structures remains unattempted. This paper seeks to address this gap by applying the DTC-VAE model to aerospace structures, aiming to achieve high HI evaluation criteria scores comparable to those achieved by semi-supervised methods. By leveraging the strengths of both Diversity-DeepSAD and DTC-VAE, the study aspires to advance HI generation for SHM in complex aeronautical environments.
Furthermore, the calculation of the typical fitness score, which includes Mo, Pr, and Tr, is prone to bias toward the training units. This bias can lead to potential issues, as the model may produce highly correlated HIs during training, resulting in misleadingly high fitness scores. However, when faced with an unmatched HI from a specific unit during testing, the performance may fall short. To address this, the rectified evaluation criteria [15, 13] are adopted in this study, ensuring a more reliable assessment during the test phase. This adjustment provides a stronger foundation for comparison and enhances the practical reliability of the standard.
To construct a historical-independent HI for aeronautical composite structures, GW-SHM emerges as a promising technique as mentioned earlier. However, each GW measurement requires data collection through a network of PZT sensors, resulting in an enormous volume of data to process. Feeding this data directly into deep learning models demands a high number of input layer nodes, which increases model complexity and computational requirements while reducing interpretability [36]. A more efficient alternative is to employ signal processing (SP) algorithms as a preprocessing step. SP algorithms often follow explicit solutions, producing globally applicable and faster outcomes. By extracting a set of statistical features from the processed signals, the size and complexity of the models can be significantly reduced. However, identifying which SP method yields the most effective features remains an open question. The proposed frameworks address this by first applying various SP algorithms to extract and compare features in terms of HIs criteria. The selected features are then fed into AI models to generate HIs, ensuring both efficiency and effectiveness in the process.
This work contributes to the field in the following ways:
  1. (i)
    Introduction of two novel historical-independent HI models designed for reliable RUL estimation with reduced data inputs.
  2. (ii)
    First application of DTC-VAE and Diversity-DeepSAD models to SHM of aerospace composite structures, specifically guided-wave SHM.
  3. (iii)
    Novel use of intermediary proxy HIs as labels in a DeepSAD-based model, moving beyond binary healthy/failed labels, enabling richer semi-supervised HI construction.
  4. (iv)
    Extension of historical-independent HI frameworks to semi- and unsupervised learning models.
The methodology employed in this study is described in detail in Section 2, with the corresponding results presented in Section 3. A discussion of these findings is provided in Section 4, culminating in the conclusions outlined in Section 5.
2 Method
This section provides a detailed overview of the methodology adopted in this study, encompassing the experimental setup and the framework for HI generation. Subsection 2.1 introduces the experimental setup which forms the foundation of the analysis. Subsection 2.2 describes the overall proposed two-stage framework for HI generation, followed by Subsection 2.3, which outlines the criteria used to evaluate the generated HIs. Subsection 2.4 delves into the signal processing techniques applied to preprocess the data, while Subsection 2.5 focuses on the feature extraction process. Finally, Subsection 2.6 presents the feature fusion approach, which include Diversity-DeepSAD, DTC-VAE, and ensemble learning models.
2.1 Experimental setup and data preparation
To evaluate the proposed framework, five single-stiffener composite panels were subjected to run-to-failure compressive fatigue loading under ReMAP project [^4] ReMAP: Real-time Condition-based Maintenance for Adaptive Aircraft Maintenance Planning. https://h2020-remap.eu/. The panels consist of IM7/8552 carbon fiber-reinforced epoxy in unidirectional prepreg, formed with layups of $\pm 45/0/45/90/-45/0]_{S}$} and $\pm 45/0/\pm 45]_{S}$} [15]. The panels were monitored using different SHM techniques [60], from which we use only GWs in order to construct HIs.
[…]
Figure 1: Single T-stiffener CFRP panel under C-C fatigue loading monitored with PZT sensors (red circles).
The input features data was arranged in 5 folds, in each of which a different specimen was used for testing while the remaining 4 were for training. Training and test data were normalised relative to the training data only using Z-score scaling, in other words, subtracting by the mean and dividing by the standard deviation.
The model output was also normalised using the training data, but with min-max normalisation, to ensure HIs had a value of approximately 0 in the first timestep (healthy state) and approximately 1 at the last (failure), although this has no effect on fitness scores.
2.2 Two-Stage framework for HI generation
This study follows a two stage framework resulting in HI generation through the use of signal processing techniques and AI models. In the two stages, as can be seen in Figure 2, time domain measurements were processed using different SP methods, including fast Fourier transform (FFT), short-time Fourier transform (STFT), empirical mode decomposition (EMD), and Hilbert transform (HT). Statistical features were then extracted from the processed signals for each path, actuator frequency, and measurement time step. As all input features should be ideally of the same SP method to ensure efficient processing, one SP method can be practically used in the workflow. With this in mind, the two methods with the highest performance were selected for further individual use.
In the third stage, as can be seen in Figure 2, the features selected were fed into AI models (fourth phase), i.e., the Diversity-DeepSAD and DTC-VAE models, where hyperparameters are fine-tuned using Bayesian optimisation. This process was repeated for each combination of excitation frequency of PZTs and SP method. The extracted HIs from each actuating frequency were then fused by an unsupervised ensemble learning model, enabling the evaluation of the remaining combinations of AI model and SP method against the HIs evaluation criteria to determine the optimal final combination.
Figure 2: Framework for HI generation and evaluation — Step 1: Data Pre-Processing, Step 2: Signal Processing, Step 3: Feature Representation, Step 4: Feature Fusion — with prognostic criteria (Mo, Pr, Tr).
2.3 HIs criteria
Given the lack of true HI labels, evaluating the quality of the developed HIs necessitates leveraging known physics-based principles and facts. The HIs criteria include Mo, Pr, and Tr, and they are standard for this application [15, 23, 61, 13, 46]. These criteria range from 0 to 1, with 1 being the ideal value.
Mo measures the extent to which the HI shows only an increasing or decreasing trend. If maintenance is not performed, health should always decrease with an increasing number of cycles, and thus a monotonic HI should more closely model reality. The modified Mann-Kendall (MMK) formulation of Mo is presented in Equation 1 [15]:
$Mo=\frac{1}{M}\sum^{M}_{m=1}\left|{\frac{1}{(N_{m}-1)}\sum^{N_{m}-1}_{i=1}\frac{\sum^{N_{m}}_{j=1,j>i}(t_{j}-t_{i})\cdot sgn(y^{m}_{j}-y^{m}_{i})}{\sum^{N_{m}}_{j=1,j>i}(t_{j}-t_{i})}}\right|$ (1)
where $M$ is the number of specimens, $N$ is the number of GW measurements (time steps), $y^{m}_{j}$ denotes a HI at timestep $j$ of specimen $m$, $t_{j}$ is the time at the $j^{th}$ measurement, $sgn(\cdot)$ is the sign function. This formulation of Mo is preferable as it considers time gaps greater than one unit, thus reducing the effect of noise and favoring monotonicity through the whole trend, as opposed to between adjacent points [15].
Pr measures the HIs deviation at failure between specimens. Preferably, all specimens should have the same value of HI at the end of life, and HIs can be reliably used to predict failure. Equation 2 expresses Pr [15], where $std(\cdot)$ refers to the standard deviation between specimens.
$Pr=exp\left(-\frac{std(y^{m}_{N})}{\frac{1}{M}\sum^{M}_{m=1}\left|{y^{m}_{1}-y^{m}_{N}}\right|}\right)$ (2)
Finally, Tr is a measure of the correlation between HIs of different units [15]. The set of specimens should show the same HI pattern from beginning to end of life, as this makes their behaviour more predictable. Tr is given by Equation 3, where $corr(\cdot)$ is the Pearson correlation function.
$Tr=min_{a,b}\;{corr(y^{a},y^{b})},\quad a,b=1,...,M$ (3)
To effectively evaluate a model’s performance on test specimens while mitigating bias towards highly matched training specimens and avoiding the obscuring of unmatched test specimen behavior, alternative evaluation criteria must be established. These criteria should prioritise the behavior of test specimens over those used during training, ensuring a more reliable assessment [15, 24]. These criteria include $Mo_{test}$, which evaluates the monotonicity of only the test specimen(s), and $Pr_{test}$, which assesses the prognosability by emphasizing the deviation of the test specimen(s) relative to those used for training. These metrics are defined in Equation 4 and Equation 5 [15]:
$Mo_{test}=\left|{\frac{1}{(N_{test}-1)}\sum^{N_{test}-1}_{i=1}\frac{\sum^{N_{test}}_{j=1,j>i}(t_{j}-t_{i})\cdot sgn(y^{m}_{j}-y^{m}_{i})}{\sum^{N_{test}}_{j=1,j>i}(t_{j}-t_{i})}}\right|$ (4)
$Pr_{test}=exp\left(-\frac{\left|{y^{test}_{N}-\frac{1}{M^{train}}\sum^{M^{train}}_{m=1}y^{m}_{N}}\right|}{\frac{1}{M}\sum^{M}_{j=1}\left|y_{1}^{j}-y_{N}^{j}\right|}\right)$ (5)
where a superscript traintrain or testtest represents train or test specimens, respectively, and thus $M^{train}$ refers to the number of the training specimens.
Finally, the fitness score $F_{all}$ is defined by Equation 6, and measures the overall quality of HIs [15], while $F_{test}$ is defined similarly in Equation 7 with a focus on test units. Setting the control constants $k_{Mo}$, $k_{Pr}$, and $k_{Tr}$ equal to 1, in order to weight the criteria equally, the maximum fitness score is 3; the fitness may also be given as a percentage relative to this maximum.
$F_{all}=k_{Mo}\cdot Mo+k_{Pr}\cdot Pr+k_{Tr}\cdot Tr$ (6)
$F_{test}=k_{Mo}\cdot Mo_{test}+k_{Pr}\cdot Pr_{test}+k_{Tr}\cdot Tr$ (7)
Both fitness scores $F_{all}$ and $F_{test}$ are calculated and discussed in this study.
2.4 Signal processing
Signal processing (SP) techniques were applied to transform the measured signals into the frequency and time-frequency domains, in order to highlight relevant information, such as frequency components or noise isolation. These transformations enabled the extraction of relevant features for further analysis. The techniques used include FFT, STFT, EMD, and HT.
The FFT algorithm transforms the original temporal signal into the frequency domain by processing the entire time domain at once and outputting corresponding amplitudes for the underlying frequencies. However, it does not retain temporal variations in the signal, such as how these frequencies evolve over time, which may reveal additional information about the structure’s health. In contrast, STFT preserves time-frequency information by dividing the data into smaller time intervals and analyzing localised segments, making it advantageous for non-stationary signals where frequency content changes over time. The result of STFT is a 2D matrix, where time and frequency are the axes, and amplitude is the value at each position. Nevertheless, there is a nonlinear trade-off between time and frequency resolution: too many time segments reduce the available frequency samples, and vice versa. In this study, it was found that performing STFT with an FFT length of 250 maximised the number of data points, resulting in an overlap length of 125.
In the case of EMD, the sifting algorithm was implemented to iteratively decompose the signal into intrinsic mode functions (IMFs) in the time domain. The process involves constructing upper and lower envelopes of the signal using cubic splines to interpolate local maxima and minima, respectively. The mean of these envelopes is then subtracted from the signal at each iteration, refining the signal until it satisfies the conditions for an IMF. Suitable stopping conditions are necessary to avoid valuable information being lost [62]. Through trial and error, these criteria were set to a minimum standard deviation of $0.1$ and a maximum of 10 iterations.
The HT algorithm produces a complex-valued signal, the analytic signal, from a real signal. The real component represents the original signal, while the imaginary component is created by phase-shifting each frequency component by $\frac{\pi}{2}$ [63]. The analytic signal therefore contains information concerning instantaneous amplitude and phase.
By employing these SP techniques, features can be extracted that capture essential signal characteristics across the time, frequency, and time-frequency domains, forming a robust foundation for subsequent analysis.
2.5 Feature extraction
Statistical features, as detailed in A, were extracted from the time, frequency, and time-frequency domains to characterise the signals comprehensively. Common statistical features for the time and frequency domains were implemented following [23].
From the time domain, 19 statistical features were extracted, as shown in Table 12. These features were computed for the original time-domain signals as well as the processed signals using HT and EMD. From the frequency domain, 14 statistical features were extracted from the FFT-transformed signals, as listed in Table 13. For the time-frequency domain, the output of STFT was divided into 17 time windows, each with discrete frequency components. To reduce dimensionality, four statistical measures—mean, standard deviation, skewness, and kurtosis—were calculated across the frequency axis within each time window. This resulted in a total of 68 features (4 measures × 17 time windows), as summarised in Table 14.
The extracted features are outlined in Table 1. By systematically extracting statistical features across multiple domains, the framework ensures a concise yet rich representation of the signal information.
Table 1: Feature numbers extracted from the results of each SP method.
Method Raw Data FFT HT EMD STFT
Features domain Time Frequency Time Time Time-Frequency
Features 1-19 20-33 34-52 53-71 72-139
For the first stage, features across different actuator-sensor paths and frequencies are averaged, resulting in 139 features per time step (GW measurement) for each specimen. After obtaining all time steps and specimens, the HI prognostic criteria defined in Subsection 2.3 are applied to evaluate each feature. To avoid anomalously poor features disproportionately impacting the SP method selected, the mean fitness score across all features is calculated to serve as a benchmark. Features with fitness scores exceeding this benchmark are retained for further analysis. The SP method yielding features with the highest remaining mean fitness scores is identified as optimal for developing an efficient framework to extract HIs.
2.6 Feature fusion
Given the extracted and selected features as inputs, two deep learning models were developed to generate HIs, i.e., semi-supervised Diversity-DeepSAD and unsupervised DTC-VAE, which will be presented in this section.
2.6.1 Diversity-DeepSAD
The original DeepSAD model incorporates labeled training data in the loss function of the otherwise unsupervised deep support vector data description (Deep SVDD) model [42]. This approach transforms input data into a hyperspace of reduced dimensions, in which the distance to a point, the hypersphere centre $\mathbf{c}$, represents the extent to which the datapoint is an anomaly. Frusque et al. [46] modified the loss function further to include a term to increase diversity, resulting in Diversity-DeepSAD for HI generation. This was demonstrated on a rotational machinery dataset with hundreds of measurement timesteps, a condition not often met in many other applications. For example, the ReMAP dataset may contain as few as 30 timesteps before failure. Consequently, it is necessary to explore the impact of alternative functions for generating auxiliary artificial labels and to evaluate the influence of the number of labeled measurements.
In prior studies [42, 46], a binary set of auxiliary labels $\tilde{y}_{j}\in\{1,-1\}$ was employed, where a label of 1 was assigned to measurements representing a very healthy condition, and -1 corresponded to those reflecting a severely damaged state. This approach focuses solely on two extreme conditions. However, these two phases typically occupy only a small portion of the system’s entire lifetime, limiting the potential to fully leverage auxiliary labels. This issue is particularly pronounced in the severely damaged phase, where measurements are rare, often with only a single observation available at the point of failure for assigning the label -1. To address these limitations, this study generated continuous auxiliary labels $\tilde{y}_{j}\subset[1,-1]$ according to Equation 8:
$\tilde{y}_{j}=1-2\frac{t_{j}}{t_{N}}$ (8)
where, for simplicity, a linear function was applied for the first and last quarters of the lifetime. The quarter ratio was established experimentally to produce improved results. As visualised in 3(a), the remaining data points do not leverage auxiliary labels, providing flexibility to the model, given the uncertainty of the true health status for these intermediate conditions.
It should be noted that auxiliary labels $\tilde{y}_{j}$ differ from the expected health indicator (HI), denoted by $y_{j}$. Assuming the hypersphere center $\mathbf{c}$ represents the healthy state, the HI is defined as the distance from that center:
[…]
where $z_{j}$ denotes the latent space representation obtained from the autoencoder. When $z_{j}$ is close to the healthy state $\mathbf{c}$, the $y_{j}$ approaches 0, indicating minimal deviation from the healthy condition. Conversely, as $z_{j}$ moves farther from $\mathbf{c}$, the $y_{j}$ increases, reflecting a more degraded condition. In this framework, the maximum value of 1 is assigned to $y_{j}$ in the fully damaged state. Consequently, the HI can freely exhibit linear or non-linear behavior—with different main trends, such as exponential or polynomial behavior—between these bounds, as illustrated in 3(b).
(a) Auxiliary labels generated for training as a function of lifetime. Data collected in the dashed region is left unlabelled.
(b) Examples of possible HI embeddings, where each curve represents a basic function capturing the main trend, either individually or in combination. Note that 0 corresponds to healthy and 1 to failure, unlike auxiliary labels (1 = healthy, -1 = failure).
Figure 3: Relationship between new auxiliary labels and embedding output expected for the proposed Diversity-DeepSAD.
Denoting the encoder and decoder functions by $\phi$ and $\psi$, respectively, the HI can be expressed as:
[…]
where $\phi_{\theta}(\mathbf{x}_{j})$ represents the latent space embedding of the input data $\mathbf{x}_{j}$ at timestep $j$, and $\theta$ denotes the learnable parameters of the encoder. The sum of squares of these embeddings, raised to the power of the auxiliary label $\tilde{y}_{j}$ where applicable, is minimised using the following loss function:
$\mathcal{L}_{DS}=\frac{1}{n+u}\sum_{i=1}^{u}\left(y_{i}\right)^{2}+\frac{\eta}{n+u}\sum_{j=1}^{n}\left(y_{j}+\epsilon\right)^{2\tilde{y}_{j}}$ (11)
in which $n$ represents the number of labeled samples, $u$ is the number of unlabeled samples, and $\eta$ is a hyperparameter that controls the contribution of the auxiliary labels. To prevent numerical zero errors when $\tilde{y}_{j}$ is negative, a small term $\epsilon$ is added to labeled HIs. This loss component is the first term in the overall Diversity-DeepSAD loss function:
$\mathcal{L}_{Diversity-DeepSAD}=\mathcal{L}_{DS}+\nu\mathcal{L}_{L2}+\lambda\mathcal{L}_{D}$ (12)
where $\nu$ and $\lambda$ are hyperparameters controlling the weights of the regularization terms $\mathcal{L}_{L2}$ and $\mathcal{L}_{D}$, respectively. The regularization terms are defined as follows:
[…]
$\mathcal{L}_{D}=\sum_{h=1}^{H}(\zeta_{h}-\ln(\zeta_{h}))$ (14)
where $L$ denotes the number of layers in the model, and $\zeta_{h}$ represents the $h^{\text{th}}$ eigenvalue of the Gram matrix $G=Z^{T}Z$, in which $Z$ is the DeepSAD embedding. $\mathcal{L}_{L2}$ represents $L_{2}$ regularization, minimizing the sum of squares of node weights to encourage a greater distribution of smaller weights and prevent overfitting, while $\mathcal{L}_{D}$ promotes a greater distribution of training data in the model output [46]. These three terms therefore work together to ensure that the model effectively learns the latent representations, while the regularization terms help prevent overfitting and promote diversity within the learned embeddings.
A standard autoencoder (AE) is always used to pretrain the network weights (learnable parameters $\theta$). This AE is formed by adding a decoder $\psi$ which inverts the Diversity-DeepSAD network layers, considering the mean squared difference between input and output data as a loss function, $\mathcal{L}_{AE}$. This loss function is used to pre-train the autoencoder, but is disregarded in the training of Diversity-DeepSAD in Equation 12. The hypersphere center $\mathbf{c}$ is initialised by computing the average output from the model during a forward pass on the training data. Initially, $\mathbf{c}$ is set to a zero matrix (e.g., dimensions $4\times 4$). In the pretraining phase, denoted by $\phi_{\theta}^{[0]}$, the model performs a forward pass on the training data, and the outputs for each sample are accumulated. The number of samples is incremented with each data point, and the cumulative outputs are added to $\mathbf{c}$. After processing all the data, the matrix $\mathbf{c}$ is normalised by the total number of samples to obtain the average. To avoid numerical issues, any component of $\mathbf{c}$ with an absolute value smaller than $\epsilon$ is adjusted as follows:
[…]
This adjustment ensures stability in the hypersphere center values, helping the model establish a reference point for the healthy state. The hypersphere center $\mathbf{c}$ serves as a critical foundation for anomaly detection and health indicator estimation. The model architecture is illustrated in Figure 4, where red, yellow and green represent pretraining, training and model outputs respectively.
Figure 4: Architecture of Diversity-DeepSAD. Dash-dotted lines indicate the pretraining autoencoder and dashed lines indicate the loss function.
Diversity-DeepSAD architecture and hyperparameter optimisation: The Diversity-DeepSAD model was implemented with 6 hidden layers, each containing half the number of neurons as the previous layer, and leaky rectified linear units (Leaky ReLU) used as activation functions between layers. The hyperspace was set to be 16-dimensional (resulting in $\mathbf{c}$ with dimensions $4\times 4$), which was found to enhance the performance of AE during pretraining. Key hyperparameters, including the number of of epochs, learning rate, and batch size—those observed to have a strong impact on the results—were optimised using Bayesian optimisation with Gaussian processes, based on the search space shown in Table 2. Maximizing the $F_{all}$ score has been set as the objective function of the Bayesian optimization algorithm. To mitigate the model’s tendency to separate the failure state too distinctly from intermediate health degradation, a low maximum learning rate was set. All remaining hyperparameters were fixed, with their values determined from literature [46], namely $\nu=10$, $\eta=10$, $\lambda=0.001$ and $\epsilon=1\cdot 10^{-6}$.
Table 2: Diversity-DeepSAD hyperparameters space.
Hyperparameter Min value Max value
Batch Size 50 150
Learning Rate (Pretraining) 0.0001 0.001
Learning Rate 0.0001 0.001
Number of Epochs (Pretraining) 5 20
Number of Epochs 50 200
2.6.2 DTC-VAE
VAEs are probabilistic generative models that encode an input vector $x_{j}$ into a lower-dimensional latent space $z_{j}$. While traditional autoencoders map $x_{j}$ to $z_{j}$ directly, VAEs instead learn a probabilistic mapping $\phi(z_{j}|x_{j})$ in the encoder. This is parametrised by a mean $\mu$ and variance $\sigma^{2}$, allowing the model to represent uncertainty in the latent space. The latent variable, which in this paper is considered as the HI, is sampled according to Equation 16, where $\varepsilon$ is Gaussian noise added for variability.
$z=\mu+\sigma\odot\varepsilon,\quad\varepsilon\sim\mathcal{N}(0,1)$ (16)
Finally, the decoder produces a reconstruction of the input $\hat{x}_{j}$ from the latent variable $z_{j}$, using the decoder probability distribution $\psi(x_{j}|z_{j})$. This mechanism is illustrated in Figure 5, where yellow and green represent training and model outputs, respectively.
Figure 5: Architecture of degradation-trend-constrained variational autoencoder (DTC-VAE).
The DTC-VAE model is optimised by minimising the loss function, $\mathcal{L}_{DTC-VAE}$, which is a summation of the Kullback-Leibler divergence ($\mathcal{L}_{KL}$), reconstruction loss ($\mathcal{L}_{RE}$) and monotonicity constraint loss ($\mathcal{L}_{Mo}$):
$\mathcal{L}_{DTC-VAE}=\alpha\mathcal{L}_{KL}(\phi,\psi)+\beta\mathcal{L}_{RE}(\phi,\psi)+\gamma\mathcal{L}_{Mo}(z)$ (17)
where $\alpha$, $\beta$, and $\gamma$ are their respective hyperparameters.
Firstly, $\mathcal{L}_{KL}$ quantifies the difference between the prior distribution and the posterior distribution approximated by the encoder, $\phi(z_{j}|x_{j})$. In this case, the prior is assumed to be a simple standard normal distribution $p(z)\sim N(0,1)$, simplifying $\mathcal{L}_{KL}$ to the form presented in Equation 18:
$\mathcal{L}_{KL}(\phi,\psi)=-\frac{1}{2}(1+\log(\sigma^{2})-\sigma^{2}-\mu^{2})$ (18)
which only includes the standard deviation $\sigma$ and mean $\mu$ of the encoder distribution.
Secondly, $\mathcal{L}_{RE}$, portrayed in Equation 19, characterises the difference between the encoder input $\mathbf{x}_{j}^{m}$ and reconstructed decoder output $\hat{x}_{j}^{m}$ for data of a specimen $m$ and timestep $j$, ensuring that these two are as similar as possible.
$\mathcal{L}_{RE}(\phi,\psi)=\sum_{m=1}^{M}\sum_{j=1}^{N_{m}}(\mathbf{x}_{j}^{m}-\mathbf{\hat{x}}_{j}^{m})^{2}$ (19)
Finally, the latent variable is prompted to follow a monotonic trend by adding the monotonicity constraint loss, $\mathcal{L}_{Mo}$. This was proposed by Qin et al. [55], ensuring the extracted HIs followed a monotonic trend by adding a degradation-trend constraint, hence the name DTC-VAE. While the other loss terms are common for VAE architectures, this one is unique to the DTC-VAE model and has a direct effect on the fitness of the output, which is expressed as follows:
$\mathcal{L}_{Mo}(z)=\sum_{j=2}^{N}(z_{j}^{m}-z_{j-1}^{m}-r)^{2}$ (20)
where $z_{j}^{m}$ represents the latent variable for specimen $m$ at timestep $j$, while the rate of degradation $r$ is a hyperparameter ensuring that $z_{j}^{m}>\left(z_{j-1}^{m}+r\right)$, which controls the magnitude of change of a HI between timesteps. Its value was determined from the original DTC-VAE study [55], in essence, a randomly selected constant in the range $(9,10)$.
The original HIs criteria functions could not be adapted for gradient descent, given the fact that they are not differentiable and require comparison of latent outputs across specimens, inaccessible during model training. In the case of monotonicity, originally presented in Equation 1, an alternate formulation $\mathcal{L}_{Mo}$ has been implemented, which ensures high monotonicity scores and is compatible with the training process. Ideally, all prognostic criteria (Mo, Pr, Tr) shall be implemented into the loss function in future to maximise fitness.
A common issue in VAEs is posterior collapse, where the latent variables fail to encode meaningful information because the decoder learns to reconstruct inputs directly and the posterior distribution collapses to the prior. This typically occurs when the KL divergence term dominates the loss [64]. In this work, the risk of collapse is mitigated through optimisation of the loss weights $\alpha$, $\beta$, and $\gamma$, following Qin et al. [55], and by the inclusion of the monotonicity constraint $\mathcal{L}_{Mo}$, which forces the latent variables to capture degradation trends instead of defaulting to the prior.
DTC-VAE architecture and hyperparameter optimisation: The DTC-VAE model consists of 3 components: the encoder, latent space and decoder. The encoder’s only hidden layer uses a sigmoid activation function and has a varying number of neurons, kept as a hyperparameter. This is followed by two output layers of size 1, representing the mean and log variance of the latent space. As explained by Equation 16, the reparameterization trick is used to allow for gradient calculation and thus backpropagation. The decoder acts as a mirror to the encoder, it takes the latent variable as input to a hidden layer with sigmoid activation functions. This hidden layer has the same number of neurons as in the encoder. It is followed by an output layer the same size as the input with linear activation function, which returns the reconstructed input. All weights are initialised with uniform Glorot (Xavier) initialisation [65] due to its state-of-the-art performance with sigmoid activation functions and improved stability. Adam is used as the optimiser.
The hyperparameters were optimised over the space detailed in Table 3, using the Bayesian optimisation algorithm maximizing the $F_{all}$ score as for Diversity-DeepSAD. From the preprocessed data and the optimised hyperparameters, a model was trained for every fold and frequency.
Table 3: DTC-VAE hyperparameters space.
Hyperparameter Min value Max value
Hidden Layer Size 40 60
Batch Size 75 95
Learning Rate 0.001 0.01
Number of Epochs 500 600
$\alpha$ 1.4 1.8
$\beta$ 2.6 3
$\gamma$ 0.05 0.1
2.6.3 Ensemble learning model
To ensure the reproducibility of results, it is essential to initialise the models with predetermined random seeds. After hyperparameter optimisation with a given seed, the models were each trained for 5 different random seeds to evaluate the stability of the models. The HIs generated by each model were then mean-averaged, and the HIs criteria applied to obtain the fitness scores reported in Section 3.
The weighted averaging ensemble (WAE) model was implemented to fuse the HI results of the different excitation GW frequencies. This is done to improve performance by reducing instability, and removing the need to select and rely on only one frequency. It was chosen to use the normalised fitness scores of the averaged HIs as weights $\tilde{\omega}_{f}$ for the same frequency, as this method was most successfully implemented previously [15]. This method is described by Equation 21, with the constants $\omega_{f}$ adjusted depending on their fitness, producing a new weighted-average HI value $\bar{y}$.
$\bar{y}=\sum_{f=1}^{F}\tilde{\omega}_{f}\bar{y}^{f}$ (21a)
$\tilde{\omega}_{f}=\frac{\omega_{f}}{\Sigma_{f=1}^{F}\omega_{f}}$ (21b)
$\omega_{f}=F_{all}\left(\bar{y}^{f}\right)$ (21c)
3 Results
In this section, signal processing methods are first compared based on the HIs criteria calculated from the statistical features extracted for each method. Following this, the results of the semi-supervised Diversity-DeepSAD model and the unsupervised DTC-VAE model are presented and compared using the HIs criteria across all specimens $F_{all}$ (training, validation, and test) as well as for test specimens alone $F_{test}$. Beyond comparing the two deep learning models, the analysis also examines the impact of different excitation GW frequencies, the fusion of all excitation GW frequencies, and the selection of signal processing methods (focusing on the two top-performing ones).
3.1 Signal Processing
Figure 6 shows the $F_{all}$ scores averaged across all paths and frequencies for each of the extracted statistical features. It is clear that all signal processing methods contain some features that perform poorly when evaluated for fitness. In order to discard the low-performing features, the set is reduced to only those performing above average ($\mu=1.614$), which is indicated by the red horizontal line in Figure 6. The mean score of features extracted from each SP method is presented in Table 4. It can be seen that after the feature reduction, FFT and HT yield the highest-performing features, which will be used for further individual use in the following AI models.
Figure 6: Fitness scores of statistical features extracted from the time domain and other domains derived from various signal processing methods.
Table 4: Mean fitness scores for each signal processing method, calculated before and after feature reduction.
Time Domain FFT HT EMD STFT
Mean $F_{all}$ before reduction 1.673 (± 0.45) 1.924 (± 0.41) 1.840 (± 0.39) 1.474 (± 0.42) 1.511 (± 0.37)
Mean $F_{all}$ after reduction 2.022 (± 0.18) 2.149 (± 0.21) 2.046 (± 0.20) 1.880 (± 0.11) 1.910 (± 0.19)
3.2 Diversity-DeepSAD
The Diversity DeepSAD model was trained across all combinations of frequencies and data folds (i.e., training and test divisions), with WAE implemented as described. The HIs constructed by Diversity DeepSAD upon FFT and HT features are shown in Figure 7 and Figure 8, respectively. For better comparison, the HIs are displayed over the normalised lifetime of specimens, ranging from 0% to 100%. The resulting fitness scores of the HIs are reported in Table 5 and Table 6. These tables include $F_{all}$ scores, from Equation 6, and $F_{test}$ scores, from Equation 7, for HIs generated from both FFT and HT features.
Table 5: Fitness scores across all units ($F_{all}$) for the Diversity-DeepSAD model using FFT and HT features over 5 iterations.
f [kHz] Fold 1 Fold 2 Fold 3 Fold 4 Fold 5
FFT HT FFT HT FFT HT FFT HT FFT HT
50 2.55 (± 0.04) 2.29 (± 0.28) 2.47 (± 0.03) 2.49 (± 0.03) 2.40 (± 0.05) 1.85 (± 0.06) 2.50 (± 0.06) 2.47 (± 0.07) 2.47 (± 0.06) 2.67 (± 0.08)
100 2.34 (± 0.07) 2.30 (± 0.05) 2.22 (± 0.05) 2.19 (± 0.14) 2.36 (± 0.04) 2.31 (± 0.04) 1.92 (± 0.19) 2.22 (± 0.10) 2.47 (± 0.03) 1.99 (± 0.11)
125 2.43 (± 0.06) 2.02 (± 0.22) 2.20 (± 0.12) 2.42 (± 0.04) 2.38 (± 0.03) 2.37 (± 0.08) 2.42 (± 0.10) 2.07 (± 0.23) 2.47 (± 0.05) 2.18 (± 0.33)
150 2.38 (± 0.12) 2.56 (± 0.11) 2.41 (± 0.04) 2.42 (± 0.06) 2.16 (± 0.14) 2.28 (± 0.08) 2.06 (± 0.11) 2.37 (± 0.14) 2.49 (± 0.06) 2.06 (± 0.14)
200 2.39 (± 0.05) 2.35 (± 0.04) 2.32 (± 0.03) 2.04 (± 0.17) 2.34 (± 0.06) 1.82 (± 0.09) 2.51 (± 0.05) 2.33 (± 0.13) 2.32 (± 0.11) 2.51 (± 0.03)
250 2.36 (± 0.05) 2.26 (± 0.04) 2.30 (± 0.05) 1.95 (± 0.13) 2.32 (± 0.05) 1.92 (± 0.32) 2.14 (± 0.07) 1.67 (± 0.19) 2.28 (± 0.10) 2.25 (± 0.15)
Fusion 2.50 (± 0.03) 2.51 (± 0.02) 2.42 (± 0.03) 2.37 (± 0.07) 2.44 (± 0.02) 2.01 (± 0.08) 2.38 (± 0.05) 2.22 (± 0.12) 2.51 (± 0.05) 2.44 (± 0.06)
Table 6: Test fitness scores ($F_{test}$) for the Diversity-DeepSAD model using FFT and HT features over 5 iterations.
f [kHz] Fold 1 Fold 2 Fold 3 Fold 4 Fold 5
FFT HT FFT HT FFT HT FFT HT FFT HT
50 2.63 (± 0.06) 2.25 (± 0.24) 2.44 (± 0.04) 2.48 (± 0.02) 2.30 (± 0.08) 1.57 (± 0.09) 2.41 (± 0.08) 2.48 (± 0.07) 2.39 (± 0.13) 2.53 (± 0.10)
100 2.37 (± 0.08) 2.26 (± 0.05) 2.17 (± 0.12) 2.00 (± 0.15) 2.21 (± 0.06) 2.12 (± 0.06) 1.67 (± 0.22) 2.13 (± 0.11) 2.40 (± 0.08) 1.64 (± 0.10)
125 2.42 (± 0.12) 1.99 (± 0.18) 2.14 (± 0.17) 2.28 (± 0.08) 2.35 (± 0.07) 2.43 (± 0.11) 2.30 (± 0.15) 2.03 (± 0.30) 2.36 (± 0.05) 2.21 (± 0.27)
150 2.29 (± 0.11) 2.58 (± 0.14) 2.42 (± 0.04) 2.27 (± 0.10) 2.31 (± 0.10) 2.32 (± 0.06) 1.75 (± 0.08) 2.30 (± 0.17) 2.35 (± 0.07) 1.93 (± 0.12)
200 2.40 (± 0.05) 2.31 (± 0.08) 2.28 (± 0.06) 2.02 (± 0.23) 2.31 (± 0.05) 1.63 (± 0.11) 2.46 (± 0.04) 2.38 (± 0.15) 2.07 (± 0.25) 2.47 (± 0.03)
250 2.36 (± 0.08) 2.18 (± 0.11) 2.17 (± 0.06) 1.59 (± 0.15) 2.23 (± 0.13) 1.77 (± 0.44) 1.88 (± 0.11) 1.56 (± 0.13) 2.17 (± 0.12) 2.15 (± 0.20)
Fusion 2.56 (± 0.03) 2.50 (± 0.03) 2.44 (± 0.04) 2.31 (± 0.07) 2.35 (± 0.05) 1.74 (± 0.06) 2.25 (± 0.08) 2.20 (± 0.09) 2.37 (± 0.09) 2.18 (± 0.09)
Figure 7: HIs constructed by Diversity-DeepSAD using FFT features.
Figure 8: HIs constructed by Diversity-DeepSAD using HT features.
Diversity-DeepSAD results show high fitness scores for both FFT and HT, with mean $F_{test}$ scores of 2.27 and 2.13 respectively, with $F_{all}$ scores averaging slightly higher at 2.35 and 2.22. This implies a general higher performance of features from the FFT than those from the HT. No fold performed consistently worst than the others, although average results from test Sample 1 were consistently highest. Inspecting the graph, this may be due to the invariably conforming behaviour of this sample, aiding in high $F_{test}$ scores. All frequencies performed similarly, with $50\text{\,}\mathrm{kHz}$ producing marginally better average $F_{all}$ results for both signal processing methods.
While the range of scores is similar for both sets of input data, FFT demonstrates a greater robustness due to its low standard deviation between random seeds, with an $F_{test}$ mean $\bar{\sigma}=$0.10 compared to 0.14 for HT. This indicates a stronger robustness in this model using the FFT features than using those of the HT.
3.3 DTC-VAE
The DTC-VAE model was similarly trained and tested on all combinations of frequencies and folds for 5 random seeds. The HIs extracted by DTC-VAE from both FFT and HT features are displayed in Figure 9 and Figure 10, respectively. The resulting fitness scores of the HIs generated by the DTC-VAE are thus provided in Table 7 and 8. These tables cover both FFT and HT features, detailing scores for the general fitness $F_{all}$ and test fitness $F_{test}$.
FFT features achieved consistently high results, with a mean $F_{test}$ scores of 2.55 and 2.20 for the FFT and HT respectively, and similar $F_{all}$ scores of 2.56 and 2.25. These imply features from the FFT generate again generally better performing HIs than the HT. No fold produced consistently different scores to the others; inspection of the graphs reveals distributed anomalies for all. However, for DTC-VAE, HIs generated using features from the $100\text{\,}\mathrm{kHz}$ signal tended to perform poorly, with an average $F_{test}$ score of 2.06 compared to the overall mean score of 2.38.
FFT features exhibited higher stability than those of the HT, with an overall lowest $F_{test}$ score of 2.24 compared to 1.42. This is further highlighted by the low standard deviation values, with an $F_{test}$ mean $\bar{\sigma}=$0.09 compared to 0.12 for HT. The stronger robustness in both models using the FFT features than those of the HT indicates that the former include more reliable information on the structure’s health, consistent with the results in Table 4.
Table 7: Fitness scores across all units ($F_{all}$) for the DTC-VAE model using FFT and HT features over 5 iterations.
f [kHz] Fold 1 Fold 2 Fold 3 Fold 4 Fold 5
FFT HT FFT HT FFT HT FFT HT FFT HT
50 2.66 (± 0.03) 2.16 (± 0.11) 2.45 (± 0.05) 2.55 (± 0.07) 2.57 (± 0.04) 2.42 (± 0.06) 2.52 (± 0.11) 2.40 (± 0.06) 2.50 (± 0.05) 2.09 (± 0.14)
100 2.37 (± 0.11) 1.96 (± 0.11) 2.41 (± 0.19) 1.88 (± 0.13) 2.55 (± 0.09) 1.71 (± 0.19) 2.52 (± 0.14) 1.61 (± 0.20) 2.45 (± 0.08) 1.50 (± 0.15)
125 2.43 (± 0.07) 2.00 (± 0.13) 2.49 (± 0.11) 2.11 (± 0.11) 2.54 (± 0.07) 2.20 (± 0.13) 2.57 (± 0.13) 2.30 (± 0.10) 2.57 (± 0.05) 2.27 (± 0.04)
150 2.69 (± 0.05) 2.63 (± 0.10) 2.54 (± 0.09) 2.59 (± 0.05) 2.53 (± 0.11) 2.60 (± 0.07) 2.51 (± 0.08) 2.62 (± 0.03) 2.75 (± 0.02) 2.65 (± 0.05)
200 2.81 (± 0.03) 2.50 (± 0.07) 2.55 (± 0.10) 2.51 (± 0.02) 2.73 (± 0.06) 2.62 (± 0.08) 2.77 (± 0.03) 2.64 (± 0.07) 2.79 (± 0.04) 2.52 (± 0.08)
250 2.35 (± 0.11) 2.06 (± 0.06) 2.66 (± 0.05) 2.19 (± 0.07) 2.44 (± 0.21) 1.97 (± 0.04) 2.51 (± 0.13) 1.99 (± 0.25) 2.67 (± 0.06) 2.16 (± 0.05)
Fusion 2.75 (± 0.01) 2.55 (± 0.02) 2.72 (± 0.05) 2.51 (± 0.07) 2.77 (± 0.07) 2.56 (± 0.05) 2.80 (± 0.03) 2.61 (± 0.04) 2.80 (± 0.03) 2.49 (± 0.02)
Table 8: Test fitness scores ($F_{test}$) for the DTC-VAE model using FFT and HT features over 5 iterations.
f [kHz] Fold 1 Fold 2 Fold 3 Fold 4 Fold 5
FFT HT FFT HT FFT HT FFT HT FFT HT
50 2.68 (± 0.02) 2.20 (± 0.13) 2.41 (± 0.07) 2.50 (± 0.08) 2.57 (± 0.07) 2.35 (± 0.10) 2.52 (± 0.08) 2.39 (± 0.08) 2.53 (± 0.04) 2.03 (± 0.15)
100 2.34 (± 0.21) 1.92 (± 0.08) 2.36 (± 0.15) 1.98 (± 0.11) 2.49 (± 0.13) 1.42 (± 0.30) 2.41 (± 0.21) 1.54 (± 0.19) 2.52 (± 0.07) 1.61 (± 0.21)
125 2.43 (± 0.06) 1.86 (± 0.19) 2.43 (± 0.14) 2.00 (± 0.13) 2.58 (± 0.08) 2.32 (± 0.08) 2.65 (± 0.10) 2.22 (± 0.14) 2.62 (± 0.08) 2.27 (± 0.06)
150 2.72 (± 0.05) 2.66 (± 0.12) 2.35 (± 0.16) 2.44 (± 0.12) 2.58 (± 0.11) 2.65 (± 0.05) 2.51 (± 0.08) 2.53 (± 0.05) 2.75 (± 0.04) 2.62 (± 0.06)
200 2.83 (± 0.02) 2.58 (± 0.09) 2.51 (± 0.15) 2.26 (± 0.10) 2.73 (± 0.06) 2.57 (± 0.07) 2.78 (± 0.02) 2.65 (± 0.07) 2.76 (± 0.05) 2.61 (± 0.05)
250 2.24 (± 0.07) 1.90 (± 0.11) 2.60 (± 0.11) 2.13 (± 0.08) 2.44 (± 0.27) 1.55 (± 0.06) 2.58 (± 0.12) 2.05 (± 0.33) 2.65 (± 0.05) 2.35 (± 0.09)
Fusion 2.76 (± 0.02) 2.60 (± 0.03) 2.71 (± 0.05) 2.43 (± 0.09) 2.76 (± 0.08) 2.41 (± 0.08) 2.77 (± 0.02) 2.61 (± 0.04) 2.81 (± 0.03) 2.61 (± 0.03)
Figure 9: HIs constructed by DTC-VAE using FFT features.
Figure 10: HIs constructed by DTC-VAE using HT features.
3.4 Weighted Average Ensemble
Overall, the fused models show improved results compared to single frequencies, with generally higher scores consistently outperforming the corresponding average fitness score of a particular fold. Across both FFT and HT, the average unfused $F_{all}$ and $F_{test}$ scores for Diversity-DeepSAD are 2.29 and 2.20, while for DTC-VAE they are 2.41 and 2.38 respectively. The same fused scores average 2.38, 2.29, 2.66 and 2.65. This is an average score increase of 3.2% for Diversity-DeepSAD, but 10.9% for DTC-VAE, with both $F_{all}$ and $F_{test}$ scores showing similar improvement. The high performance of WAE, fusing different GW excitation frequencies, combined with DTC-VAE is also evident from the graphs, where the fused HIs consistently behave more cleanly than their constituent frequencies.
The stability of fused results is also greatly improved, with deviations between folds smaller than within individual frequencies. For both models, but still most noticeably DTC-VAE, the fused graphs are visually more consistent between folds than unfused. Between seeds, the standard deviations of fitness for the fusion are always below 0.1, with one exception in Diversity-DeepSAD fold 4. The average standard deviation for Diversity-DeepSAD-Fusion scores is half that of unfused, 0.05 compared to 0.10 for $F_{all}$ and 0.06 to 0.12 for $F_{test}$. For DTC-VAE, the reduction is better than half at 0.04 from 0.09 and 0.05 from 0.11 respectively. These low standard deviations indicate that HIs fused using WAE of multiple guided wave excitation frequencies are more robust. This is because different GW excitation frequencies can capture various damage modes in structures, and the ensemble approach reduces susceptibility to errors in individual constituent models.
Across all results, there are only three occurrences where the standard deviation of fused results for HIs from the FFT exceeds that of the HT, confirming that the features extracted from the output of FFT also tend to produce more consistently performing HIs. The $F_{test}$ scores for Diversity-DeepSAD with WAE average 2.39 and 2.19 for FFT and HT respectively, while DTC-VAE shows a large increase to 2.76 and 2.53. Every DTC-VAE WAE fitness score is higher than its Diversity-DeepSAD counterpart, showing exceptionally high performance at a peak of 2.81 or 94%.
3.5 Model comparison
The prognostic criteria and fitness scores (reported in the previous subsections and tables) were averaged across all folds for both models to comprehensively perform a comparison, including the WAE models. The resulting averages for $F_{all}$, along with the prognostic criteria scores, are displayed in Table 9, while those for $F_{test}$ are included in Table 10.
Table 9: Prognostic criteria and fitness scores ($F_{all}$) averaged across all folds.
Score Diversity-DeepSAD Diversity-DeepSAD WAE DTC-VAE DTC-VAE WAE
FFT HT FFT HT FFT HT FFT HT
$\boldsymbol{Mo}$ 0.87 (±0.04) 0.87 (±0.06) 0.94 (±0.01) 0.93 (±0.03) 0.92 (±0.06) 0.85 (±0.07) 0.99 (±0.01) 0.93 (±0.03)
$\boldsymbol{Pr}$ 0.89 (±0.07) 0.85 (±0.16) 0.91 (±0.03) 0.84 (±0.18) 0.97 (±0.04) 0.87 (±0.13) 0.98 (±0.01) 0.90 (±0.02)
$\boldsymbol{Tr}$ 0.59 (±0.13) 0.50 (±0.17) 0.60 (±0.08) 0.54 (±0.11) 0.68 (±0.13) 0.52 (±0.22) 0.80 (±0.05) 0.71 (±0.05)
$\boldsymbol{F_{all}}$ 2.35 (±0.16) 2.22 (±0.28) 2.45 (±0.06) 2.31 (±0.20) 2.56 (±0.16) 2.25 (±0.34) 2.77 (±0.05) 2.54 (±0.06)
Table 10: Test prognostic criteria and test fitness scores ($F_{test}$) averaged across all folds.
Score Diversity-DeepSAD Diversity-DeepSAD WAE DTC-VAE DTC-VAE WAE
FFT HT FFT HT FFT HT FFT HT
$\boldsymbol{Mo_{test}}$ 0.83 (±0.04) 0.82 (±0.12) 0.93 (±0.04) 0.91 (±0.05) 0.91 (±0.10) 0.81 (±0.13) 0.98 (±0.02) 0.91 (±0.10)
$\boldsymbol{Pr_{test}}$ 0.85 (±0.13) 0.81 (±0.24) 0.86 (±0.10) 0.74 (±0.29) 0.97 (±0.05) 0.88 (±0.16) 0.97 (±0.02) 0.91 (±0.09)
$\boldsymbol{Tr}$ 0.59 (±0.13) 0.50 (±0.17) 0.60 (±0.08) 0.54 (±0.11) 0.68 (±0.13) 0.52 (±0.22) 0.80 (±0.05) 0.71 (±0.05)
$\boldsymbol{F_{test}}$ 2.27 (±0.23) 2.13 (±0.34) 2.39 (±0.12) 2.19 (±0.26) 2.55 (±0.18) 2.20 (±0.38) 2.76 (±0.06) 2.53 (±0.11)
Both models showed promising monotonicity scores, with Diversity-DeepSAD achieving an average $Mo$ score of 0.87 and 0.87 for FFT and HT features respectively, while its average $Mo_{test}$ scores were 0.83 and 0.82. DTC-VAE achieved average $Mo$ scores of 0.92 and 0.85 and average $Mo_{test}$ scores 0.91 and 0.81.
The average prognosability scores were also high, particularly for DTC-VAE. Diversity-DeepSAD achieved an average $Pr$ score of 0.89 and 0.85 for FFT and HT features respectively. Its average $Pr_{test}$ scores were 0.85 and 0.81. On the other hand, DTC-VAE achieved average $Pr$ scores of 0.97 and 0.87, demonstrating exceptional performance on FFT features. Similarly, its average $Pr_{test}$ scores were 0.91 and 0.81.
The results reflect the fact that trendability was the lowest performing prognostic criterion for both models, as can be seen by the inconsistency of shapes between specimens in the HI graphs. This could be due to significant uncertainties introduced during the manufacturing and fatigue experiments, as explained in Subsection 2.1. It shows that the developed framework satisfies the first two fact-based criteria while striving to fulfill the aspirational trendability metric to capture the overall trend. Both models achieved moderate scores, with Diversity-DeepSAD scoring an average $Tr$ score of 0.59 and 0.50, and DTC-VAE scoring 0.68 and 0.52, for FFT and HT features respectively.
Overall, HIs extracted by Diversity-DeepSAD using FFT features performed slightly better than using HT features. Similarly, DTC-VAE also generated better HIs when trained on FFT features, although the difference in performance was greater. However, both models showed significant improvements when combined with WAE, with all prognostic criteria showing improved scores and stability for both FFT and HT features. Diversity-DeepSAD WAE achieved average $F_{all}$ scores of 2.45 and 2.31, while its $F_{test}$ scores were 2.39 and 2.19. In addition, DTC-VAE WAE scored 2.77 and 2.54 in $F_{all}$, and 2.76 and 2.53 in $F_{test}$.
4 Discussion
By analysing Diversity-DeepSAD and DTC-VAE, key differences in their performances and behaviors are revealed. These differences shed light on their respective strengths and limitations in extracting meaningful HIs for composite structure prognostics.
Diversity-DeepSAD graphs show a tendency for HIs to increase in sharp steps, which could indicate different damage states during the fatigue life of composite structures. However, these increases were sometimes erroneous or abrupt; this is most visible in Figure 8 in test Sample 3 at $50\text{\,}\mathrm{kHz}$ and test Sample 4 at $125\text{\,}\mathrm{kHz}$, showing the model may be sensitive to sudden changes in test sample structural health or sensor damage. On the other hand, DTC-VAE did not demonstrate the same behaviour, showing a smoother increase over the specimen lifetime, indicating improved performance in this regard. This is particularly the case with FFT features, with some exceptions in the case of HT features.
Furthermore, results from both DTC-VAE and in particular Diversity-DeepSAD showed a tendency for HIs to spike at EoL, such as test Sample 3 at $125\text{\,}\mathrm{kHz}$. This is due to the model learning the failure state as a separate category to the continuous range of degradation up until that point, an effect which can be emphasised with alternative hyperparameters. This suggests that the deviation in the structural health condition after failure is significantly greater and distinct from the condition before failure, demonstrating the damage accumulation process in composite structures. Although high prognosability scores may indicate that extracted HIs are more suitable for RUL prediction, it is crucial to ensure that changes in HI values are not too abrupt. For example, a sharp increase near the end of life may leave insufficient time to take maintenance action. In this regard, DTC-VAE could potentially provide HIs that are more useful than Diversity-DeepSAD.
The improved performance of DTC-VAE can be linked to the connection between its loss function and the fitness of the extracted HIs. The incorporation of the monotonicity constraint into its loss function makes part of the fitness metric differentiable, and thus, optimisable. This provides a direct relationship between training and output fitness, by penalizing the model for producing HIs with lower fitness scores, ensuring that these follow a smooth, monotonic trend. This trend is easier to predict, making the extracted HIs more suitable for prognostics.
The incremental changes in HIs produced by Diversity-DeepSAD could on the one hand be seen as a more interpretable result, potentially illustrating the distinct damage states of composite structures. This could provide valuable insights for prognostic models, especially state-based ones, in predicting RUL. However, the increases in HIs should not be too abrupt, and the overall trend must meet the criteria; otherwise, using HIs for RUL prediction would be challenging, as discussed previously. Further studies and testing are needed to establish a stronger connection between these phases and actual physical damage states. On the other hand, DTC-VAE resulted in overall smoother HIs, with slightly higher fitness scores and less sharp jumps, that could improve RUL prediction.
Based on the average prognostic criteria and fitness scores across all units, Diversity-DeepSAD and DTC-VAE demonstrate competitive performance compared to other proposed models, as shown in Table 11. In this table, two SHM techniques—guided wave (GW) and acoustic emission (AE)—are also compared. It should be noted that the constructed HIs using AE are inherently time-dependent due to the nature of the sensing technology, whereas the GW technique has the potential to provide data for designing historical-independent HIs, as demonstrated by the proposed models in this study and HT-SSCNN [15].
While SSLSTM using AE data [23] achieves the highest fitness score across all units ($F_{all}$) with a performance of 93%, it relies on test units for validation during the LOOCV process to stop training and fine-tune hyperparameters. This reliance potentially limits its applicability to real-world scenarios where test units are new and unknown. However, this limitation has been addressed in subsequent work [24], where CEEMDAN-SSLSTM followed by a BiLSTM ensemble achieved the best performance for AE data, with 91.3% for $F_{all}$ and 86.3% for test units ($F_{test}$).
On the other hand, using GW data, the proposed historical-independent HIs in this study result in the lowest deviation between fitness scores obtained for all units ($F_{all}$) and test units ($F_{test}$), demonstrating reduced overfitting. Both the semi-supervised Diversity-DeepSAD model and the unsupervised DTC-VAE model are significantly more stable, with much lower standard deviations compared to existing state-of-the-art models. The DTC-VAE model, followed by WAE fusion, achieved the highest performance of 92% for test units ($F_{test}$) and 92.3% for all units ($F_{all}$). Notably, the deviation between fitness scores for all units and test units is minimal (0.3%).
Table 11: Performance of various methods in the ReMAP dataset collected using guided wave (GW) and acoustic emission (AE) techniques.
Prognostic
Criteria
Diversity-DeepSAD
on FFT (WAE)
DTC-VAE
on FFT (WAE)
HT-SSCNN [15]
SSLSTM [23]
(simple ensemble)
CEEMDAN-SSLSTM [24]
(BiLSTM ensemble)
$\boldsymbol{Mo}$ 0.94 (±0.01) 0.99 (±0.01) - 0.99 (±0.01) -
$\boldsymbol{Pr}$ 0.91 (±0.03) 0.98 (±0.01) - 0.86 (±0.14) -
$\boldsymbol{Tr}$ 0.60 (±0.08) 0.80 (±0.05) - 0.93 (±0.03) -
$F_{all}$ 2.45 (±0.06) 2.77 (±0.05) 2.78 (±0.15) 2.79 (±0.14) 2.74 (±0.19)
$\boldsymbol{Mo_{test}}$ 0.93 (±0.04) 0.98 (±0.02) - - -
$\boldsymbol{Pr_{test}}$ 0.86 (±0.10) 0.97 (±0.02) - - -
$\boldsymbol{Tr}$ 0.60 (±0.08) 0.80 (±0.05) - - -
$F_{test}$ 2.39 (±0.12) 2.76 (±0.06) 2.67 (±0.20) - 2.59 (±0.24)
Number of units 5 composite specimens 12 composite specimens
SHM technique GW AE
5 Conclusion
This study proposed a novel plenary data-driven framework for constructing health indicators (HIs) for aeronautical composite structures by developing two advanced machine learning models: Diversity-DeepSAD and DTC-VAE. First, statistical features from the best-performing signal processing techniques, fast Fourier transform (FFT) and Hilbert transform (HT), were extracted as inputs for the AI models. This step enabled dimensionality reduction of guided wave signals, each 2000 data points in length, recorded by PZTs across 56 actuator-sensor inspection paths at six excitation frequencies (2000×56×6 data points per measurement), to just a few tens of features per excitation frequency. This allows for simpler deep learning architectures while preserving critical information. Feature selection was based on fitness scores, calculated as the sum of the prognostic criteria of monotonicity, prognosability, and trendability. FFT features demonstrated superior performance, achieving a performance of 71.0%, i.e., an average fitness score of 2.13 (out of 3.00), compared to 68.2% for HT features.
Diversity-DeepSAD and DTC-VAE exhibited distinct behaviors in HI construction. Diversity-DeepSAD produced incremental changes in HIs, potentially reflecting distinct damage states, but occasionally showed abrupt increases, particularly near the end of life (EoL). In contrast, DTC-VAE generated smoother HIs with higher fitness scores, benefiting from its monotonicity-constrained loss function, which directly optimized the fitness metric. This smoother trend makes DTC-VAE more suitable for remaining useful life (RUL) prediction, as it avoids abrupt changes that could limit actionable maintenance time. Therefore, the augmented Diversity-DeepSAD is recommended as a closer representative of true health state, while DTC-VAE may be more useful for prognostic applications.
The proposed framework outperformed state-of-the-art models in constructing historical-independent HIs. DTC-VAE, followed by weighted averaging ensemble fusion, achieved the highest test fitness score of 2.76 (92% performance) with minimal deviation (0.3%) from the fitness score across all units, indicating reduced overfitting and improved generalization. Moreover, the models demonstrate greater stability in performance compared to existing approaches. Unlike some existing models, which rely on test units for validation, the proposed framework ensures applicability to real-world scenarios where test units are unknown.
Overall, this study provides an efficient and robust framework for HI construction, leveraging the strengths of Diversity-DeepSAD and DTC-VAE. These models enhance the accuracy and stability of HIs while supporting more reliable RUL predictions, contributing to improved maintenance strategies and operational safety for aeronautical composite structures. Future work will focus on extending the DTC-VAE loss function to incorporate additional prognostic criteria and validating the framework on diverse datasets to further enhance its applicability.
Appendix A List of features by domain
Table 12: Time domain features [23].
No Feature Equation No Feature Equation 1 Mean value $X_{m}=\frac{\Sigma^{N}_{i=1}x(i)}{N}$ 11 Shape factor $X_{shape}=\frac{X_{rms}}{\frac{1}{N}\Sigma^{N}_{i=1}\lvert x(i)\rvert}$ 2 Standard deviation $X_{sd}=\sqrt{\frac{\Sigma^{N}_{i=1}(x(i)-X_{m})^{2}}{N-1}}$ 12 Impulse factor $X_{impulse}=\frac{X_{peak}}{\frac{1}{N}\Sigma^{N}_{i=1}\lvert x(i)\rvert}$ 3 Root amplitude $X_{root}=\left(\frac{\Sigma^{N}_{i=1}\sqrt{\lvert x(i)\rvert}}{N}\right)^{2}$ 13 Max-min difference $X_{p2p}=max\left(x(i)\right)$     $min\left(x(i)\right)$ 4 Root mean square $X_{rms}=\sqrt{\frac{\Sigma^{N}_{i=1}(x(i))^{2}}{N}}$ 14 -17 $k$th central moment ($k=3,4,5,6$) $X_{k\_cm}=\frac{\Sigma^{N}_{i=1}\left(x(i)-X_{m}\right)^{k}}{N}$ 5 Residual sum of squares $X_{rss}=\sqrt{\Sigma_{i=1}^{N}\lvert x(i)\rvert^{2}}$ 6 Peak maximum $X_{peak}=max(|x(i)|)$ 18 FM4 $X_{FM4}=\frac{X_{4\_cm}}{X_{sd}^{4}}$ 7 Skewness $X_{skewness}=\frac{\Sigma^{N}_{i=1}\left(x(i)-X_{m}\right)^{3}}{(N-1)X_{sd}^{3}}$ 19 Median $X_{med}=\frac{\Sigma^{N}_{i=1}t(i)}{N}$ 8 Kurtosis $X_{kurtosis}=\frac{\Sigma^{N}_{i=1}\left(x(i)-X_{m}\right)^{4}}{(N-1)X_{sd}^{4}}$ 9 Crest factor $X_{crest}=\frac{X_{peak}}{X_{rms}}$ 10 Clearance factor $X_{clearance}=\frac{X_{peak}}{X_{root}}$
Table 13: Frequency domain features [23].
No Feature Equation No Feature Equation 1 Mean frequency $S_{1}=X_{mf}=\frac{\Sigma^{N_{f}}_{f=1}s(f)}{N_{f}}$ 8 - $S_{8}=\sqrt{\frac{\Sigma^{N_{f}}_{f=1}f_{k}^{4}s(f)}{\Sigma^{N_{f}}_{f=1}f_{k}^{2}s(f)}}$ 2 (same as variance) $S_{2}=\frac{\Sigma^{N_{f}}_{f=1}\left(s(f)-S_{1}\right)^{2}}{N_{f}-1}$ 9 - $S_{9}=\frac{\Sigma_{f=1}^{N_{f}}f_{k}^{2}s(f)}{\sqrt{\Sigma_{f=1}^{N_{f}}s(f)\Sigma_{f=1}^{N_{f}}f_{k}^{4}}}$ 3 (same as skewness) $S_{3}=\frac{\Sigma^{N_{f}}_{f=1}\left(s(f)-S_{1}\right)^{3}}{N_{f}\left(\sqrt{S_{2}}\right)^{3}}$ 10 - $S_{10}=\frac{S_{6}}{S_{5}}$ 4 (same as kurtosis) $S_{4}=\frac{\Sigma^{N_{f}}_{f=1}\left(s(f)-S_{1}\right)^{4}}{N_{f}S_{2}^{2}}$ 11 - $S_{11}=\frac{\Sigma_{f=1}^{N_{f}}(f_{k}-S_{5})^{3}s(f)}{N_{f}S_{6}^{3}}$ 5 - $S_{5}=X_{fc}=\frac{\Sigma_{f=1}^{N_{f}}f_{k}s(f)}{\Sigma_{f=1}^{N_{f}}s(f)}$ 12 - $S_{12}=\frac{\Sigma_{f=1}^{N_{f}}(f_{k}-S_{5})^{4}s(f)}{N_{f}S_{6}^{4}}$ 6 - $S_{6}=\sqrt{\frac{\Sigma_{f=1}^{N_{f}}(f_{k}-S_{5})^{2}s(f)}{N_{f}}}$ 13 - $S_{13}=\frac{\Sigma_{f=1}^{N_{f}}\sqrt{(f_{k}-S_{5})}s(f)}{N_{f}\sqrt{S_{6}}}$ 7 - $S_{7}=X_{rmsf}=\sqrt{\frac{\Sigma_{f=1}^{N_{f}}f_{k}^{2}s(f)}{\Sigma_{f=1}^{N_{f}}s(f)}}$ 14 - $S_{14}=\sqrt{\frac{\Sigma_{f=1}^{N_{f}}(f_{k}-S_{5})^{2}s(f)}{\Sigma_{f=1}^{N_{f}}s(f)}}$
Table 14: Time-frequency domain features [66].
No Feature Equation
1 Mean value $X_{m}=\frac{\Sigma^{N}_{i=1}x(i)}{N}$
2 Standard deviation $X_{sd}=\sqrt{\frac{\Sigma^{N}_{i=1}(x(i)-X_{m})^{2}}{N-1}}$
3 Skewness $X_{skewness}=\frac{\Sigma^{N}_{i=1}\left(x(i)-X_{m}\right)^{3}}{(N-1)X_{sd}^{3}}$
4 Kurtosis $X_{kurtosis}=\frac{\Sigma^{N}_{i=1}\left(x(i)-X_{m}\right)^{4}}{(N-1)X_{sd}^{4}}$